I need help solving my equation using elimnation method.Solving for both x andy
Problem:
2x+y=-1
-x+2y=13
I have tried several different ways in order to find the answer, and have asked my parents for help if you could help me I would be very very very thankful

Answer by elima(1433)

2x+y=-1
-x+2y=13
2x+y=-1
y=-1-2x
Now plug this into the eq. to solve for x;
-x+2(-1-2x)=13
-x-2-4x=13
-5x=13+2
-5x=15
x=-3
Now that you havex, solve for y;
2(-3)+y=-1
-6+y=-1
y=5
Hope you understand
=)